# Hamiltonian actions on 0-shifted cosymplectic groupoids

### v1

@@ 2. Precosymplectic Hamiltonian actions
Let us consider the toric symplectic manifold $(\mathbb{C}^{n},\mathbb{T}^{n},\sum_{j=1}^{n}dz_{j}\wedge d\bar{z}_{j})$ with moment map $\mu:\mathbb{C}^{n}\to(\mathbb{R}^{n})^{\ast}\cong\mathbb{R}^{n}$ defined by $\tilde{\mu}(z_{1},\ldots,z_{n})=\sum_{j=1}^{n}(|z_{j}|^{2}-1)\xi_{j}$. Here $\{\xi_{j}\}$ stands for the canonical basis of $\mathbb{R}^{n}$. If $\mathfrak{n}=\langle\xi_{1},\ldots,\xi_{k}\rangle$ then
@@
    Let $\phi:(G^{\prime}\rightrightarrows M^{\prime})\to(G\rightrightarrows M)$ be a Morita map. Then, the cochain map $\flat_{(\omega,\eta)}$ is a quasi-isomorphism at all points in $M$ if and only if the cochain map $\flat_{(\phi^{\ast}\omega,\phi^{\ast}\eta)}$ is so at all points in $M^{\prime}$. In particular, the pullback $\phi^{\ast}:\Omega_{\textnormal{bas}}^{\bullet}(G)\to\Omega_{\textnormal{bas}}^{\bullet}(G^{\prime})$ gives rise to a one-to-one correspondence between 0-shifted cosymplectic structures on $G$ and $G^{\prime}$.
@@
These facts can be proven by arguing exactly as in Lemma 3.2.2 and Propositions 3.2.3 and 3.2.4 from [23] since the differential forms we are dealing with here are basic. ∎

### v2

@@ 2. Precosymplectic Hamiltonian actions
Let us consider the toric symplectic manifold $(\mathbb{C}^{n},\mathbb{T}^{n},\sum_{j=1}^{n}dz_{j}\wedge d\bar{z}_{j})$ with moment map $\tilde{\mu}:\mathbb{C}^{n}\to(\mathbb{R}^{n})^{\ast}\cong\mathbb{R}^{n}$ defined by $\tilde{\mu}(z_{1},\ldots,z_{n})=\sum_{j=1}^{n}(|z_{j}|^{2}-1)\xi_{j}$. Here $\{\xi_{j}\}$ stands for the canonical basis of $\mathbb{R}^{n}$. If $\mathfrak{n}=\langle\xi_{1},\ldots,\xi_{k}\rangle$ then
@@
    Let $\phi:(G^{\prime}\rightrightarrows M^{\prime})\to(G\rightrightarrows M)$ be a Morita map [^1] It is well-known that we can consider the Morita map to cover a surjective submersion on objects.. Then, the cochain map $\flat_{(\omega,\eta)}$ is a quasi-isomorphism at all points in $M$ if and only if the cochain map $\flat_{(\phi^{\ast}\omega,\phi^{\ast}\eta)}$ is so at all points in $M^{\prime}$. In particular, the pullback $\phi^{\ast}:\Omega_{\textnormal{bas}}^{\bullet}(G)\to\Omega_{\textnormal{bas}}^{\bullet}(G^{\prime})$ gives rise to a one-to-one correspondence between 0-shifted cosymplectic structures on $G$ and $G^{\prime}$.
@@
These facts can be proven by arguing exactly as in Lemma 3.2.2 and Propositions 3.2.3 and 3.2.4 from [23] since the differential forms we are dealing with are basic. Nevertheless, here we sketch a few details about the second assertion for the reader’s convenience. Since $\phi$ is a Morita map, it follows that the chain map
[…]
is a quasi-isomorphism for all $x\in M^{\prime}$. More explicitly:
$\ker(\rho^{\prime})=\text{Lie}(G^{\prime}(x,x))\simeq\text{Lie}(G(\phi(x),\phi(x)))=\ker(\rho),$
and
$\text{coker}(\rho^{\prime})\simeq N_{x}M^{\prime}\simeq N_{\phi(x)}M\simeq\text{coker}(\rho).$
Therefore, the Lie groupoid $G^{\prime}$ is a foliation groupoid if and only if so is $G$. To verify our remaining quasi-isomorphism condition observe that
$\langle\phi^{*}\flat_{x}(u),v\rangle=\phi^{*}\flat_{x}(u,v)=\flat_{\phi(x)}(d\phi(u),d\phi(v))=\langle\flat_{\phi(x)}(d\phi(u)),d\phi(v)\rangle,$
for $u,v\in T_{x}M^{\prime}$. If we assume that $\text{im}(\rho_{\phi(x)})=\ker(\flat_{\phi(x)})$ and $u\in\ker(\phi^{*}\flat_{x})$, since $\phi$ can be taken as a submersion, then $d\phi(u)\in\ker(\flat_{\phi(x)})$ and $d\phi(u)]$ is zero in $\text{coker}(\rho)$. Hence, $[u]$ is zero in $\text{coker}(\rho^{\prime})$. Thus, $\ker(\phi^{*}\flat_{x})=\text{im}(\rho^{\prime}_{x})$. Similarly, if we suppose that $\text{im}(\rho^{\prime}_{x})=\ker(\phi^{*}\flat_{x})$ and $\tilde{u}\in\ker(\flat_{\phi(x)})$, then $\tilde{u}$] is zero in $\text{coker}(\rho)$. That is to say, $\ker(\flat_{\phi(x)})=\text{im}(\rho_{\phi(x)}).$
From the above we get that if $\flat$ defines a quasi-isomorphism for all point in $M$, then so does $\phi^{*}\flat$ for all points in $M^{\prime}$. Conversely, suppose that $\phi^{*}\flat_{x}$ defines a quasi-isomorphism for every point $x\in M^{\prime}$. In this case, given $y\in M$ we know that there exist a point $x\in M^{\prime}$ together with an arrow $y\xleftarrow{\,g\,}\phi(x)$, as the map $\phi$ is Morita. Consequently, the fact that if $\flat$ is a quasi-isomorphism at a point then it is so at every element in the orbit through such a point, allows us to guarantee that $\flat_{y}$ is a quasi-isomorphism. ∎